OpenStudy (anonymous):

OK well today I learned about the "constant of proportionality" (7th grade math). We use the equation y=kx. k= constant of proportionality. I had homework problems on that, had the same equation where we substituted with a scenario. I dont understand what k represents. Could u help by giving an ex problem (solved) so i can look at it and see how to complete it. Also could u write what k represents? Thank a ton guys

OpenStudy (rizags):

k is essentially an arbritary constant defining y/x. This is because y/x = k. In other words, y is in direct proportion to x, such that y/x = k. Thus, given k, you can find values of y and x that satisfy, or given y and k, you can find x.

OpenStudy (anonymous):

what?

OpenStudy (rizags):

ok so, k is just a number. doesn'r really matter unless it is SPECIFICALLY ddefined by the problem. for example, if i'm given, find 2 values of y and x that satisfy y=kx and k=6, those variables y and x could be 36 and 6, or 6 and 1, because 36/6 = 6 and 6/1 = 6.

OpenStudy (anonymous):

could u give a word problem then help "define" k @Rizags

OpenStudy (rizags):

Ok here, For a given spring, F has a value of 35 when the spring has stretched 8 inches. What is the constant of proportionality for that spring? so F=kx in this scenario when F=35, x=8 therefore, 35 = k(8), so k=35/8

OpenStudy (anonymous):

k represents what?

OpenStudy (anonymous):

and whats x?

OpenStudy (rizags):

the constant of force (proportionality) in the given problem. its a Constant, so as x and y vary, K stays the same.

OpenStudy (rizags):

this is kind of a physics application tho

OpenStudy (anonymous):

i meant what does k represent like as in is it for a given problem "a gallon of gas cost"

OpenStudy (rizags):

so, in a gas price problem, if y= the total price of gas, and x= the gallons u put, then K is the constant= gas price

OpenStudy (rizags):

K ALLLLWAYS stays the same

OpenStudy (anonymous):

ok thanks..

OpenStudy (rizags):

medal rl quik plz lol

OpenStudy (anonymous):